OpenStudy (anonymous):

Use the quadratic formula to solve 3x2 + 4x - 6 = 0.

OpenStudy (anonymous):

\[a ^{2}+bx+c = 0\] use this formula \[x = \frac{ -b \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

*correction \[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]